I need help turning a standard equation into slope- intercept form. Can you help me with that?
My equations are:
3x - 6y = 9
4x - y - 3 = 0
x + y = 0
x + 3y - 3= 0
y - 0.5 = 0
2x + 3y - 4 = x + 5
and help me do the last equation as well.

Answer by mananth(16946)

3x-6y=9
add -3x to both sides
-3x+3x-6y=9-3x
-6y= 9-3x
/-6
y= 9/-6 -3x/-6
y= -3/2 +(1/2)x
y= (1/2)x -3/2

2x + 3y - 4 = x + 5
add -2x to both sides
-2x+2x+3y-4 = x+5 -2x
3y-4= -x+5
add +4 to both sides
3y+4-4=-x+5+4
3y=-x+9
/3
y= -x/3 + 3